OpenStudy (anonymous):

a rabbit has walked 20 meters and a turtle has walked b centimeters from the same direction. write an expression to represent the distance that the turtle needs to walk to catch up the rabbit. a.give the answer in meters b.give the answer in centimeters needddd help anyone???

OpenStudy (anonymous):

r=20 t=b he needs to walk 20-b meters i don't really know metric system too well, so I can't really help with the cm part... but im guessing its 20-b(100) because isn't 100cm=1m?

OpenStudy (anonymous):

x=distance that the turtle needs to walk to catch up the rabbit a/. x=20-(b/100) b/. x=20*100-b